OpenStudy (anonymous):

How you solve 13 - 3/2x = 37 Help!!!

OpenStudy (tkhunny):

Add (3/2)x to both sides?

OpenStudy (anonymous):

Help!!!!

OpenStudy (tkhunny):

Did you do it?

OpenStudy (anonymous):

No, the answer is wrong The answers are -7 and -16

OpenStudy (tkhunny):

You are jumping too quickly. Two things: 1) Just add (3/2)x to both sides and show me what you get. 2) This equation, as written, has only one solution. Let's check the problem and the answer you believe to be correct, and make sure we're on the same page.

OpenStudy (anonymous):

I think the answer is 7 2/5

OpenStudy (tkhunny):

The answer can't be positive. Well, did we get an intermediate result, yet?

OpenStudy (anonymous):

So is it - 7 2/5?

OpenStudy (tkhunny):

It can't be 5th's, either. I'm still waiting for @Itz_Menor to demonstrate one step.

OpenStudy (anonymous):

I dont know how to do it

OpenStudy (tkhunny):

Did you add (3/2)x to both sides? I am asking you to do one thing. I'm not asking you to know how to do it. Just add that quantity to both sides of the equation and show me what you get.

OpenStudy (anonymous):

I got 8

OpenStudy (anonymous):

1.85 or 1 85/100

OpenStudy (anonymous):

I give up

OpenStudy (anonymous):

9- 2x = 35 -2x=35-9 x= 35-9/2 x=13

OpenStudy (tkhunny):

13 - 3/2x = 37 Add (3/2)x to both sides. REALLY!! 13 = 37 + (3/2)x Subtract 37 -24 = (3/2)x Multiply by 2/3 -16 = x You MUST show intermediate results and steps you can follow. Otherwise, there is no way to help you. "I got 8" tells me nothing.

OpenStudy (tkhunny):

...except that you invented your own problem statement.